OpenStudy (winsteria):

SOMEONE PLEASE HELP ME IF YOU DO I'LL GIVE YOU A MEDAL AND FAN What is the product in lowest terms? 5 ~8 --- ∗ --- 6 15 A) 8 ~ --- 9 B) 4 ~ --- 9 C) 1 ~ --- 3 D) 1 ~ --- 7 (IMAGINE THE ~ SIGNS AS NEGATIVE SIGNS)

OpenStudy (winsteria):

@aum @mony01 @TheSmartOne

OpenStudy (anonymous):

B

OpenStudy (anonymous):

\[(5/6) *(-8/15) = -40/90 \] = -4/9

OpenStudy (winsteria):

Thank you both so much